OpenStudy (anonymous):

Can anyone help? Square root of P = 10p

OpenStudy (anonymous):

Answer is supposed to be a solution set..

OpenStudy (anonymous):

\[ \sqrt{p}= 10 p \Rightarrow 100p^2-p =0 \Rightarrow p(100p-1)=0 \Rightarrow p = 0,\frac 1{100}\]